OpenStudy (azionne):

7(x -2) + 3(x + 2) = 5(x - 6) medal

pooja195 (pooja195):

Same concept start by distributing \[\huge~\rm~\bf{7}(x -2) + 3(x + 2) = 5(x - 6) \] \[\huge~\rm~7(x-2)\] \[\huge~\rm~7\times x=?\]

pooja195 (pooja195):

Or you can do it all at once \[~\rm~(7)(x)+(7)(−2)+(3)(x)+(3)(2)=(5)(x)+(5)(−6)\]

pooja195 (pooja195):

@Azionne are you there??

OpenStudy (azionne):

yea

pooja195 (pooja195):

.-. can you tell me what equation we would get after distributing?

OpenStudy (azionne):

7x-14-3x+6=5x-36

pooja195 (pooja195):

\[\huge~\rm~7x-14+3x+6=5x-30\] Good now combine like terms \[\huge~\rm~\bf~(7x+3x)+(−14+6)\]

OpenStudy (azionne):

21x-8

pooja195 (pooja195):

\[\huge~\rm~10x−8=5x−30\] Now to isolate x our next step is to subtract 5x from both sides \[\huge~\rm~10x-5x−8=5x-5x−30\] 10-5=?

OpenStudy (azionne):

5

pooja195 (pooja195):

5-5= ?

OpenStudy (azionne):

0

pooja195 (pooja195):

Good so now we have \[\huge~\rm~5x−8=−30\] Now we need to add 8 to both sides \[\huge~\rm~5x−8+8=−30+8\] \[\huge~\rm~\bf-30+8=?\]

OpenStudy (azionne):

-22

pooja195 (pooja195):

Good -8+8=0 <---cancelled out \[\huge~\rm~5x=−22\] Next to isolate x we need to divide both sides by 5x \[\huge~\rm~5x/5=−22/5\] What is our final answer?

pooja195 (pooja195):

5/5=?

OpenStudy (azionne):

1

pooja195 (pooja195):

Good so you get just x -22/5 cannot be simplified further therefore it remains the same Final answer- x=-22/5

rvc (rvc):

good job